According to the Heisenberg uncertainty principle, what is the minimum possible 1D box size (L) capable
of trapping an electron (me = 9.109 × 10−31 kg)
Answer #1

According to the Heisenberg uncertainty principle::

\Deltax.\Deltap = \frac{h}{4\pi } ......................(1)

where h=6.6*10^-34 j.s

  \Delta x=1mm

Me = 9.1*10^-31 kg and

find   \large \Delta v=??

\large \Deltap =Me*\large \Delta v.............(2)

we can calculate that \large \Delta x *\large \Delta p = \large \frac{h}{4\pi }   \large \approx10^-34

put the value in equation number (2)

(10^-3m)(10^-30)\large \Delta v =10^-34

10^-33m.kg(\large \Delta v)=10^-34

\large \Delta v=10^-1=0.1m/s

\large \Delta v=0.1 m/s